what is 4/5 of 35? please show me how you got the answer .
Answer by Alan3354(69443)

what is 4/5 of 35?
-------------
(4/5)*(35/1) = 4*35/5 = 140/5
= 28
Does that answer it for you?
-------------
4/5 = 0.8
0.8*35 = 28